# Frequency Is What You Need: Word-frequency Masking Benefits Vision-Language Model Pre-training

### v1

@@ Appendix C More results
More detailed results are presented in this section.
C.1 Zero-shot Classification on More Datasets
In Table 16, we present the evaluation results of CLIPF following the approach of CLIP [3] on additional datasets. In almost all of the downstream datasets, when using six text tokens, CLIPF surpasses not only the baseline CLIP without image and text masking and FLIP with image masking and full-text tokens but also outperforms other text masking strategies. Similar to observations from the ImageNet datasets, reducing the number of text tokens from 16 to 8 substantially enhances CLIPF’s performance across all datasets. Specifically, performance improvements are noted: SUN397 [38] by 6.39%, DTD [39] by 9.31%, OxfordPets [40] by 4.64%, STL10 [41] by 9.31, EuroSAT [42] by 6.60%, UFC101 [43] by 6.68% and ImageNet-1K [44] by 3.13%. The performance of other text masking strategies, including truncation, random, block, and syntax, also improves with this reduction. However, the performance gains are not as large as those observed with CLIPF. This improvement suggests that employing text masking strategies as a data augmentation technique can enhance VLMs’ performance in zero-shot classification tasks. CLIPF, utilizing a word frequency masking method, demonstrates robust zero-shot classification performance and establishes a benchmark for state-of-the-art results.
C.2 Zero-shot Image-text Retrieval
@@
We report the details of zero-shot image-text retrieval for Flickr30k and COCO datasets, the details are shown in Table 17.
@@
C.3 Learn Curved before fine-tuning
In Figure 1, we show the learning curve after fine-tuning. Moreover, we show the learning curve before fine-tuning, they have a similar performance for every text masking strategy.
@@ C.3 Learn Curved before fine-tuning
Figure 7: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning.

### v2

@@ Appendix B More Results
More detailed results are presented in this section, including the results and learn curve of image-text retrieval.
B.1 The details of Image-text Retrieval
@@
In Table 12, we show more details of Zero-shot Image-Text Retrieval on COCO and Flickr30k datasets.
@@
B.2 Learn Curved for Image-text Retrieval
We also investigate the zero-shot image-text retrieval performance over training epochs on the COCO and Flickr30k datasets. It also shows that the number of training epochs has a substantial impact on the performance, as illustrated in Fig. 9. And the performance gap between syntax masking and the other text masking on the zero-shot image-text retrieval task is larger than zero-shot classification task. This is because the zero-shot image-text retrieval task requires the model not only to understand the objects/nouns, but also to understand the logic of the text. Moreover, we present the zero-shot image-text retrieval performance across training epochs for models pre-trained on the LAION-400M dataset. As illustrated in Fig. 10, the performance patterns of models trained on LAION-400M are similar to those of previous models that pre-trained on the CC12M dataset. Specifically, CLIPF consistently outperforms syntax masking. Block masking achieves comparable performance to syntax masking and even surpasses it on the image-to-text retrieval task using the COCO dataset at epoch 16th.
Figure 9: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model for an additional epoch without image and text masking.
Figure 10: Zero-shot Image-Text Retrieval accuracy on COCO and Flick30k over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on LAION-400M for 16 epochs with 25 image tokens and 4 text tokens to speed up training and fine-tune the model for 0.4 epoch with 224x224 image size and 32 text tokens.
Additionally, Fig. 11 illustrates the validation loss across pre-training epochs. It is evident that the validation loss for truncation and syntax masking methods begins to increase earlier compared to other text masking methods, suggesting that these methods may overfit to the data sooner.
Figure 11: The validation loss curve of COCO and Flick30k datasets over the pre-training epochs.
B.3 Learn Curved for DTD and OxfordPets Datasets
The learning curves for the DTD [40] and OxfordPets [41] datasets are presented because not all class names in these datasets are nouns. As shown in Table 13, the performance of syntax masking on the DTD dataset is consistently lower than that of random, block, and frequency masking across all epochs, rather than initially outperforming them and declining at later stages in the ImageNet dataset, as shown in Fig. 12. Moreover, the performance of syntax masking is very poor and almost close to truncation masking. OxfordPets is a dataset containing incomplete noun class names. In the early training epochs, syntax masking performs similarly to frequency masking and still outperforms truncation, block, and random masking. However, in the later stages of training, syntax masking performs much worse than the other text masking strategies.
@@ B.3 Learn Curved for DTD and OxfordPets Datasets
Figure 12: Zero-shot classification accuracy on ImageNet-1K dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Figure 13: Zero-shot classification accuracy on DTD dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
Figure 14: Zero-shot classification accuracy on OxfordPets dataset over training epochs for CLIPF and CLIPA strategies. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training and fine-tune the model additional epoch without image and text masking.
B.4 Learn Curved before fine-tuning
In Fig. 1, we show the learning curve after fine-tuning. Moreover, we show the learning curve before fine-tuning in Fig. 15 (30 epochs) and Fig. 16 (45 epochs). They have a similar pattern to the models before fine-tuning for every text masking strategy.
Figure 15: Zero-shot classification accuracy on ImageNet-1K over the epochs for the different text masking strategies before fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 30 epochs with image and text masking (75%) to speed up training.
Figure 16: Zero-shot classification accuracy on ImageNet-1K over the 45 epochs for the different text masking strategies before fine-tuning. The backbone of the image encoder is ViT-B/16, and the model is pre-trained on CC12M for 45 epochs with image and text masking (75%) to speed up training.
B.5 Fine-tuning the models with More Epochs
The purpose of fine-tuning the models without image and text masking is to reduce the distribution gap between pre-training and evaluation. As demonstrated in FLIP and CLIPA, a small number of fine-tuning steps is usually sufficient to bridge this gap [8, 12]. We conducted additional experiments using two fine-tuning epochs, shown in Tab 13, to demonstrate that increasing the number of fine-tuning epochs does not substantially improve performance but instead leads to higher computational costs.
Table 13: Comparison of CLIPF, CLIP, and CLIPA for zero-shot classification on ImageNet-1K. Reported as top-1 accuracy. All models use a ViT-B/16 image encoder backbone pre-trained on CC12M for 30 epochs with 75% image masking for speedup (pre-train). Then, we fine-tune the model with one epoch (fine-tune-1) and two epochs (fine-tune-2) without image and text masking.
Models Masking Image Tokens Text Tokens CC12M
pre-train fine-tune-1 fine-tune-2
CLIP 197 32 36.6
FLIP 49 32 32.0 33.7 34.1
CLIPA truncation 49 8 25.4 28.4 28.5
random 34.5 36.9 37.1
block 35.5 37.9 38.0
syntax 28.5 35.0 35.5
CLIPF frequency 36.6 39.3 39.4
B.6 Comparison with SW-CLIP
Table 14: Comparison of the zero-shot accuracy performance of SW-CLIP and CLIPF on ImageNet-1k. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Datasets Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
SW-CLIP CC3M 49 6 14.9 16.8
CLIPF 14.4 18.2
SW-CLIP CC12M 49 8 35.9 38.5
CLIPF 36.6 39.3
B.7 Applying text masking on SigLIP
We also applied the text masking strategies to SigLIP; the results are presented in Tab. 15. CLIPF still achieves superior performance compared to other text masking strategies and without text masking. This further supports our conclusion that frequency-based masking is the more effective strategy.
Table 15: We apply text masking (75%) on SigLIP to pre-train on CC12M for 30 epochs with 75% image masking for speedup pre-training. Reporting the zero-shot classification accuracy on ImageNet-1K.
Models Masking Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
SigLIP 197 32 39.3
SigLIP 49 32 28.4 29.6
SigLIP truncation 49 8 20.4 27.1
random 29.7 31.7
block 30.5 32.7
syntax 25.9 31.6
SigLIP frequency 32.4 34.8
Appendix C Ablation
model Image tokens Text tokens Threshold ViT-B/16
pre-train fine-tune
CLIPF 49 8 1e-5 35.6 38.4
1e-6 36.6 39.3
1e-7 36.1 38.6
C.1 Threshold Analysis
Table 17: Comparison of CLIPF pre-trained with token masking and word masking on ImageNet-1K for zero-shot classification. The models are pre-trained on CC12M [15] for 30 epochs with image masking (75%) to speed up training and fine-tune the model an additional epoch without image and text masking.
Models Masking Image Tokens Text Tokens ViT-B/16
pre-train fine-tune
CLIPF frequency-token 49 8 36.0 38.0
frequency-word 49 8 36.6 39.3
C.2 Word and Token Analysis
Since Open_CLIP encodes text using byte-pair encoding (BPE) [9], some words are represented by multiple tokens. Therefore, in this experiment, we calculate the masking probability based on token frequency rather than whole words. Table 17 compares the performance of both approaches. Although token frequency masking may disrupt the original word structure and thus achieve lower performance compared to word frequency masking, it still slightly outperforms other text masking strategies.